OpenStudy (anonymous):

Use the arc length formula and the given information to find s. Show your work for full credit. r = 20 ft θ = 19° s = ?

OpenStudy (ghazi):

theta= arc/radius

OpenStudy (anonymous):

thats the formula

hero (hero):

Hint: \(s = r\theta\)

OpenStudy (ghazi):

19=arc/20...therefore arc= 19*20

hero (hero):

You have to convert theta to radians first

OpenStudy (ghazi):

yes agreed!! 19 degrees have to be converted to radians first

OpenStudy (anonymous):

how do i do that?

OpenStudy (ghazi):

1degree = pi/180 radians

OpenStudy (anonymous):

so 19 degree is pi/?

OpenStudy (ghazi):

19 degrees= pi*19/180= 0.105pi

OpenStudy (anonymous):

ohh ok so thats the final?

hero (hero):

s = arc length, so theta always has to be in radians.

OpenStudy (ghazi):

yes..now 0.105*pi radians have to be multiplied to 20 to get the arc length

OpenStudy (anonymous):

is that 2.10/pie

OpenStudy (ghazi):

yess

OpenStudy (ghazi):

2.10pi

OpenStudy (anonymous):

yay !! i got half of it !

OpenStudy (ghazi):

done :)

OpenStudy (anonymous):

whats the final answer then?